The compound 1-pentanol, C5H12O, is a good fuel. It is a liquid at ordinary temperatures. When the liquid is burned, the reaction involved is

2 C5H12O(ℓ) + 15 O2(g)10 CO2(g) + 12 H2O(g)


The standard enthalpy of formation of liquid 1-pentanol at 25 °C is -351.6 kJ mol-1; other relevant enthalpy of formation values in kJ mol-1 are:
C5H12O(g) = -294.6 ; CO2(g) = -393.5 ; H2O(g) = -241.8

(a) Calculate the enthalpy change in the burning of 5.000 mol liquid 1-pentanol to form gaseous products at 25°C. State explicitly whether the reaction is endothermic or exothermic.

ΔH° =
kJ

(b) Would more or less heat be evolved if gaseous 1-pentanol were burned under the same conditions?


What is the standard enthalpy change for vaporizing 5.000 mol C5H12O() at 25°C?

ΔH° =
kJ


Calculate the enthalpy change in the burning of 5.000 mol gaseous 1-pentanol to form gaseous products at 25°C.

ΔH° =
kJ
Answer #1

E5€ +6_+O_= 64,0(1) -351-6 kJ/mol o n C+02 - COLG) = -393-5 kJ/mol - © H+ 0. = 4,06) D#- 241-8 kg/mol - © Multiplying equatioThis Mal According to Hears Law, Alt for this Reaction is a =(b) – frox393-5)-(124240-8) = 6247-4 kJ/mol under the same Mare